OpenStudy (gamenerd123):

What is the average rate of change of the function f(x)=2(3)x from x = 2 to x = 4?

OpenStudy (sooobored):

is the problem \(f(x)=2(3)^x\)?

OpenStudy (gamenerd123):

Yes

OpenStudy (sooobored):

next time, i suggest you write the problem as f(x) = 2(3)^x instead ^ denotes exponents when x=2 what is 2(3)^x =? when x=4 what is 2(3)^x=?

OpenStudy (gamenerd123):

It would be \[f4=2(3)^{4}\], right?

OpenStudy (sooobored):

f(4) yes

OpenStudy (gamenerd123):

Just f(4)?

OpenStudy (sooobored):

f(4) = 2(3)^4 the parenthesis is important now, when you punch the right side into a calculator, what value do you get?

OpenStudy (gamenerd123):

\[f(4)=2(3)^{4}\]?

OpenStudy (sooobored):

3^4 = 81 2*81 = ?

OpenStudy (gamenerd123):

162

OpenStudy (sooobored):

right, so when x=4, then f(4) =162

OpenStudy (sooobored):

so when x=2, f(2) = ?

OpenStudy (gamenerd123):

Ok, so should I just write 162 or actually write f(4)=162?

OpenStudy (sooobored):

write f(4)=162, though thats not the answer

OpenStudy (gamenerd123):

Okay so one question, how did we get 81?

OpenStudy (sooobored):

oh, im guessing you're not familiar with powers if you have x*x*x*x*x its not fun to write out 5 x's so instead you represent it as \(x^5\)

OpenStudy (gamenerd123):

Oh right, 3x3x3x3

OpenStudy (sooobored):

so 3^4=\(=3^4=3*3*3*3=9*9=81\)

OpenStudy (gamenerd123):

Okay understood now :)

OpenStudy (sooobored):

ok, so the 2nd step, what is f(2)=? or f(x) when x=2?

OpenStudy (gamenerd123):

So when 2 is X 2=4

OpenStudy (gamenerd123):

2x2=4

OpenStudy (sooobored):

why? the function f(x) = 2(3)^x what happened to the 3?

OpenStudy (sooobored):

all you're doing is just replacing the x with 2

OpenStudy (gamenerd123):

Oh

OpenStudy (gamenerd123):

So 2 would either be f(2) or (3)^2

OpenStudy (sooobored):

no... when you substitute 2 in, you get f(2) = 2(3)^2 cant separate/ignore the equal sign

OpenStudy (sooobored):

f(2) = 2(3)^2 so what does 2(3)^2 =?

OpenStudy (gamenerd123):

9?

OpenStudy (sooobored):

3^2 = 9 so 2(3^2) = 2*9 = ?

OpenStudy (gamenerd123):

18

OpenStudy (sooobored):

ok so f(2) = 18 yes?

OpenStudy (sooobored):

i mean ify ou want to show your steps, which teachers love looking at, you can write it f(2) =2(3)^2 = 18 or just break it up into 2 steps

OpenStudy (gamenerd123):

So 18 is the answer?

OpenStudy (sooobored):

f(2) = 2(3)^2 f(2)=18

OpenStudy (gamenerd123):

And yea I would like to do that

OpenStudy (sooobored):

no... what is your question asking? copy and paste it

OpenStudy (gamenerd123):

What is the average rate of change of the function f(x)=2(3)x from x = 2 to x = 4?

OpenStudy (gamenerd123):

??

OpenStudy (sooobored):

ok, average rate of change it is the equivalent of saying , what is the slope?

OpenStudy (sooobored):

between the two points x =2 and x=4

OpenStudy (gamenerd123):

Isn't it 3?

OpenStudy (gamenerd123):

That was wrong, wasn't it?

OpenStudy (gamenerd123):

What is the avergae rate of change then?

OpenStudy (gamenerd123):

Cuz I'm just lost

OpenStudy (sooobored):

sorry, forgot i was still doing this do you remember what slope is? or what hte slope formula is ?

OpenStudy (gamenerd123):

Slope is something is basically the term that happens evently, like how many gallons he/she used per mile. Sorry I'm not exactly a good explainer

OpenStudy (sooobored):

or miles per gallon, this is a rate, similar to rate of change being asked in your problem but more generally, slope ona graph describes how steep the curve is

OpenStudy (gamenerd123):

Oh

OpenStudy (sooobored):

now, if you have 2 points, then if you remember the slope formula, \[m=\frac{y_2 - y_1}{x_2-x_1}\]

OpenStudy (sooobored):

except this case, we have no y actually, you can consider f(x) as y so f(2) = y2 and f(4) = y1, x2= 2 and x1 =4

OpenStudy (sooobored):

substitute that all in and tell me what m would be m is slope

OpenStudy (gamenerd123):

oh uhh...

OpenStudy (gamenerd123):

I'm sorry for giving a hard time, I'm not very sure

OpenStudy (gamenerd123):

My mind is just so stupid when it comes to equation formulas

OpenStudy (sooobored):

ok, that the information you already have , put it in terms of ordered pairs f(4) = 162 :: (4,162) f(2)= 18 :: ( 2, 18)

OpenStudy (gamenerd123):

Ahh okay, that makes sence

OpenStudy (sooobored):

does that make it a bit more clear? now, apply these points to the slope formula \[ m= \frac{y_1-y_2}{x_1-x_2}\]

OpenStudy (sooobored):

the resulting value would be the slope aka rate of change

OpenStudy (gamenerd123):

But how would I apply these numbers to the slope formula?

OpenStudy (sooobored):

ordered pairs come in the form (x,y)

OpenStudy (sooobored):

always

OpenStudy (gamenerd123):

Is the answer 9?

OpenStudy (gamenerd123):

I divided 162 by 18, and I just thought that it might be a way to solve it

OpenStudy (sooobored):

doesnt matter which order you do the subtraction, as long as its in the same order top and bottom that sounds a bit low \[\frac{162-18}{4-2}\]

OpenStudy (sooobored):

\[\frac{ 162-18 }{4-2}=\frac{18-162}{2-4}\]

OpenStudy (gamenerd123):

So do I subtract those?

OpenStudy (sooobored):

162- 18 =?

OpenStudy (gamenerd123):

144

OpenStudy (sooobored):

4-2=?

OpenStudy (gamenerd123):

2

OpenStudy (sooobored):

144/2=?

OpenStudy (gamenerd123):

72

OpenStudy (sooobored):

m= 72 now you understand how to find slope from 2 points?

OpenStudy (sooobored):

or rate of change between 2 point?

OpenStudy (gamenerd123):

Oh, lol math is such a Jerk XD and yes

OpenStudy (sooobored):

math has an specific order in which to do things, this order is unchanging, as long as you understand the order, things really are not that hard though i do suggest you memorize the slope formula or basic theory on how you find slope from 2 points, you will probablly see it alot in the future, so better to learn it now rather than having to constantly relearn it over and over and over and over again

OpenStudy (gamenerd123):

Okay, I've taken notes from this discussion

OpenStudy (sooobored):

also, if you;re familiar with PEMDAS or BODMAS or BEDMAS fractions if you want to write it linearly are written [numerator]/[denominator]

OpenStudy (gamenerd123):

Bodmas and Bedmas? Never heard of those

OpenStudy (sooobored):

writing x1-x2/y1-y2 = \(x_1 -\frac{x2}{y1} -y2\)

OpenStudy (sooobored):

in canada i think its Bedmas- bracket exponent etc.

OpenStudy (gamenerd123):

Oh okay

OpenStudy (sooobored):

some places call it bracket , order, and switch multiplicatio nadn division

OpenStudy (sooobored):

all the same thing though, "order of operations"

OpenStudy (gamenerd123):

lol I'll remember that

OpenStudy (sooobored):

well, good luck with the rest

OpenStudy (gamenerd123):

Thx so much, you were such a great help

OpenStudy (sooobored):

and a study tip, while this website is still up, try to help people with their problem and explain how to get the answer (since last i heard its closing next year) try to understand the concept well enough that you are able to explain it other, a bit like reviewing the concept and they might ask you questions you mightve never thought of when initially learning

OpenStudy (gamenerd123):

Will do, but I have to go right now